Question 5:
A box contains 4 bad and 6 good bulbs. Two are selected randomly. One of them is tested and found good. What is the probability that other will be good too?
“A” student in a class can solve 75% of the problems and another student “B” can solve 70%. What is the probability either A or B can solve a problem if chosen at random?

Question 6:
Let A and B are two events such that P(A)= ¼, P(A/B)=1/2, P(B/A)=2/3
Let A and B are independent events
A & B are mutually exclusive events.
Find P(A∩B) and P(B)

Question 7:
Two coins are tossed. What is the conditional probability that two heads result, given that there is at least one head?

Question 8:
Two fair dice, one red and one green, are thrown. Let A denote the event that the red die shows an even number and B, the event the green die shows 5 or 6. Show that event A & B are independent.
[hint: you may take help from last file of probability handouts there mentioned formulas of dependent/independent events]
1
Expert's answer
2020-05-11T11:29:21-0400

Question 5: 

Sample space for two bulbs

"(good, good), (good, bad), (bad, good), (bad, bad)"


"P(good, good)={\dbinom{6}{2}\dbinom{4}{0} \over \dbinom{10}{2}}={1 \over 3}"

"P(good)={6 \over 10}={3 \over 5}"

"P(good|good)={P(good, good) \over P(good)}={{1 \over 3} \over {3 \over 5}}={5 \over 9}"

"P(A)=0.75, P(B)=0.7"

"P(A^C)=1-P(A)=1-0.75=0.25"

"P(B^C)=1-P(B)=1-0.7=0.3"

"P(A^C\cap B^C)=P(A^C)\cdot P(B^C)=0.25\cdot 0.3=0.075"

"P(A\cup B)=1-P(A^C\cap B^C)=1-0.075=0.925"

Question 6:

Let A and B are two events such that P(A)= ¼, P(A/B)=1/2, P(B/A)=2/3

Let A and B are independent events

A & B are mutually exclusive events.

Find P(A∩B) and P(B) 


"P(A|B)={P(A\cap B) \over P(B)}, P(B|A)={P(A\cap B) \over P(A)}""P(A)=1\/4, P(A|B)=1\/2, P(B|A)=2\/3"

"P(A\cap B)=P(A)P(B|A)={1 \over 4}\cdot{2 \over 3}={1 \over 6}"

"P(B)={P(A\cap B) \over P(A|B)}={ {1 \over 6}\over {1 \over 2}}={1 \over 3}"

Independent events: "P(A\cap B)=P(A)P(B)."


"P(A|B)=P(A), P(B|A)=P(B)"

Mutually exclusive events: "P(A\cap B)=0"


"P(A|B)=P(B|A)=0"

Question 7: 

Sample space for two coins


"(HH), (HT), (TH), (TT)"

"P(HH|at\ least \ H)={P(HH) \over P(at\ least \ H)}={ {1 \over 4}\over {3 \over 4}}={1 \over 3}"

Question 8:

Sample space for two dice

"(R1, G1), (R1, G2), (R1, G3), (R1, G4), (R1, G5), (R1, G6),"

"(R2, G1), (R2, G2), (R2, G3), (R2, G4), (R2, G5), (R2, G6),"

"(R3, G1), (R3, G2), (R3, G3), (R3, G4), (R3, G5), (R3, G6),"

"(R4, G1), (R4, G2), (R4, G3), (R4, G4), (R4, G5), (R4, G6),"

"(R5, G1), (R5, G2), (R5, G3), (R5, G4), (R5, G5), (R5, G6),"

"(R6, G1), (R6, G2), (R6, G3), (R6, G4), (R6, G5), (R6, G6),"



"P(A)={18 \over 36}={1 \over 2}"

"P(B)={12 \over 36}={1 \over 3}"

"P(A\cap B)={6 \over 36}={1 \over 6}={1 \over 2}\cdot {1 \over 3}=P(A)P(B)"

The events "A" and "B" are independent.